Startup Night 9 - 20170910 UT

Log transferred from: https://docs.google.com/document/d/1QATuV1rrwet_xMmdMHlxGyHrXPpXnWjAIkf6TEGsJBw/

Tucson: Jenny, Dave, John Hill

Off Site: Olga

Mountain: Geno, John Morris, James, Ilya

Summary: Thick clouds, Rain, High Humidity to start, clearing just after 1am. Closed dome time used for flexure verification and FS/pupil alignment. Seeing too poor for AGw transform data collection, so passed on to Pepsi where progress was made guiding on-axis with both AGW3&4. LBC skyflats in the morning.

Night priorities outlined: https://docs.google.com/spreadsheets/d/1q3hQ04UESRPDJTzQn5SQ1EOXBOee_conayY7WJYgpRQ/edit#gid=0



19:45 James swapped back in Controller 1 into LBCR and Olga taking test biases

21:04 LUCI flexure verification. Olga is running biases at different rotator positions, working in simultaneously.

luci#.20170910.#

  1. Imaging - Optic Sieve N375 5 sec short delay at 90deg

    1. Luci1 0005-25

    2. Luci2 0009-29

  2. Imaging - Optic Sieve N375 20sec delay @ 78deg

    1. Luci1 0026-46

    2. Luci2 0030-50

  3. Imaging - Optic Sieve N375 20sec delay @ 35deg

    1. Luci1 0047-67

    2. Luci2 0051-71

  4. Spec - spectrosieve N180 G210 20sec delay @35deg

    1. Luci1 0069-89

    2. Luci2 0072-92

  5. Spec - spectrosieve N180 G210 20sec delay @78deg

    1. Luci1 0090-110

    2. Luci2 0093-113

  6. DARKS

    1. LUCI1 DIT 5sec 111-115

    2. LUCI2 DIT 5sec 114-118

@~21:00 LUCI Observer GUI on rm525 computer froze. Took luci_snapshot. Decided to run through x2go session on obs2 instead.

@21:28 GEIRS Error 1 Command \x91sync save\x92 returned error code 1. Frame buffer is empty (not yes read) on LUCI1. Did a read through LUCI1 GEIRS, and was able to resume.

@22:50 another GEIRS sync error for LUCI2 this time. Just a read of geirs2 solved it

@2300 another geirs sync error for LUCI2. THis time FITS error persisted so restarted GEIRS2 services through LMC

22:58 Olga is noting that the banding for LBCR appears when the filterwheel is moving during readout. She continues investigations.

07:36 - 09:42 a series of 96 biases taken with no movement of rotator of filterwheel all look flat. No banding. Consistent with the series of almost 300 taken the previous night. T301 ~ 13.5 C for all of these on both nights.

LUCI2 Field stop & Pupil alignment

J N30 Image 0123 center 1026.65, 1037.38

Pupil image 0124 973.79, 1111.82 -> plug 993.978, 1104.10

dXimg dYimg dXpup dYpup

-2.65 -13.38 -20.18 7.72

FM1m1 FM1m2 FM4m1 FM4m2

-93 -2603 -2036 2836

LUCI2 verification images pupil 0125, FS 0126

LUCI1 Field stop & Pupil alignment

J N30 Image 0117 center 1033.44, 1025.61

Pupil image 0118 1127.32, 1116.61,-> plug 1058.42, 1106.68

For these N180 camera pixel offsets:

dXimg dYimg dXpup dYpup

-9.44 -1.61 68.89 9.93

Motor offsets to align both the Pupil and FieldStop are:

FM1m1 FM1m2 FM4m1 FM4m2

1489 7762 224 -2818

Verification images pupil 0119, FS 0120

1:20 Opening

1:35 pointing check

1:40 Seeing 1.1\x94

2:08 Flat was not loaded on DX. Geno loaded the preshutdown test.201307 flat and we had a star!

Unsuccessful attempt at LUCI1 Transform Data

2:11 Authorizing LUCI1 only

2:16 LUCI1 AGw Transform data collection Stone_B1 luci1.20170910.0120-

2:18 Seeing blew up and cloud forming over target. Seeing 2.2-2.5 on guider. 0.8-1.0 on dimm?. Data crap

2:44 Seeing showing no signs of improving. Handing over to Pepsi for their guide transform work. Geno reconfiguring.

PEPSIPOL Guiding Configuration On-Axis

05:30 ACQUIRE preset has a possible star at +x and -x

OSA391: gs=7 gives x=85.4 y=72.1 (native 85.4,540)

DX side AGw4 guides OK at RIGHTZEROPOINT=-135 - probably SFP_rotation isn't used by GCS

06:29 SX needs guidingCorr_flip_x changed to -1.0. Are the two guider CCDs different?

06:39 BS9116 - SX needs pointingCorr_flip_x changed to -1.0

Re: Pepsi Pol Work, Quoting John

\x93The success of tonight was that we are now able to guide on-axis with AGW3 and AGW4 . This took much longer than I expected because:

a) There was a CCD clipping configuration error on the guider images related to a change in GCS ~1 year ago.

b) I found that adjusting GCS SFP_rotation had no effect, but also was not needed for guiding.

c) AGW3 needed an X-flip on the guider while AGW4 did not. I'm still puzzling about why that is the case.

The failure of tonight was that we were not able to locate off-axis guide stars. That makes me worry that the AGW probe is on the wrong side of the field, but it could be something much more simple like a 5 degree error in the ZEROPOINT of the field rotation."

LBC Sky Flats

11:46 Switching to LBC\x92s for skyflats

12:08 Sky Flats LBCB: B, V LBCR: R, z

Lbcb.20170910.120815 -122200 B @PA0

Lbcb.20170910.122445-122750 B @PA180

Lbcb.20170910.123135-123420 V@PA 180

lbcr.20170910.121816-122155.fits R @PA 0

lbcr.20170910.122442-122746.fits R @PA 180

lbcr.20170910.123130-123416.fits z @PA 180

12:28 too bright to get the Vz 0deg. Closing up

[From the archive, it looks like Us Y flats were taken at the end of the night, too. Was

It this set that was too bright?] Jenny: This set was saturated. Definitely too bright. Sorry excluded from log

Taking 10 lbc biases

-- %USERSIG{BarryRothberg - 2017-09-10}%

Comments

 
Topic revision: r4 - 26 Oct 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback